ADNAN MUHAMMED

Grade 12,

p { margin-bottom: 0.25cm; direction: ltr; line-height: 120%; text-align: left; }p.western { } A gravitational force of acts on an object and displaces it from to 4 . What is the work done by the gravitational force?

p { margin-bottom: 0.25cm; direction: ltr; line-height: 120%; text-align: left; }p.western { }
  1. A gravitational force of acts on an object and displaces it from to 4 . What is the work done by the gravitational force?

Grade:11

1 Answers

Nishant Vora IIT Patna
7 years ago
F = -20j
dr = (4i -6j) – (2i+6j ) = (2i – 12 j)

Now word done is F.dr = 20*12 = 240

Thanks